OpenStudy (anonymous):

As you stop your car at a traffic light, a pebble becomes wedged between the tire tread. When you start off, the distance of the pebble from the pavement in relation to the distance you have traveled can be modeled with a sinusoidal function. Assume that the diameter of the wheel is 24 inches. Which equation BEST models this situation?

OpenStudy (anonymous):

A) y = 12cos( 1 12 x) + 12 B) y = -12cos( 1 12 x) + 24 C) y = -12cos( 1 12 x) + 12 D) y = -12cos( 1 12 x) - 12 Eliminate E) y = -12cos([[24π]]x) + 12